In a card game (using a standard deck of cards), you pay $5 to draw a single card. If you draw an Ace, you win $20. If you draw a face card, you win $10. If you draw a 2, you win $5.. . a) Fill in the probability distribution below:. . X51020. P(X) . . . b) What is the expected value?. c) Is the game fair? Why or why not?
p$5 = 4/52 =1/13 p$10 = 16/52 =3/13 p$20 = 4/52 =1/13 now expected value of x is the sum of x* P(x), which = 5/13 + 30/13 + 20/13 = 55/13 hope it helps